# Manifolds, Random Matrices and Spectral Gaps: The geometric phases of generative diffusion

### v1

@@ A.2 Double variance scenario
Once the density of the eigenvalues is computed one can perform the same change of variables described in Appendix A.1 for the case of single variance, and obtain the density $\rho_{t}(r)$ for the eigenvalues of $J_{t}$.
Figures (4) and (4) report the evolution in time of the spectral density, as well as its cumulative function, when $f=0.25$ and $f=0.75$. The cumulative function has been used to estimate the formation of the intermediate gaps to compare with the experiments for the estimation of the data-manifold dimension.
@@ Appendix B Network training and Model Architecture Details
For the image datasets, we used the diffusion setting in (Ho et al.,, 2020). We use the variance scheduler with $\beta_{\text{min}}=10^{-4}$ and $\beta_{\text{min}}=2\times 10^{-2}$, $T=1000$ time steps, and score model backbone (PixelCNN++ (Salimans et al.,, 2017)). Furthermore, for each of the datasets, we adjusted the partameters to account for the different complexity see Table 1.
For the linear models, we used a Variance Exploding continuous score model trained with 2M steps (batch size 128). The model had a Residual architecture with size 128 hidden channels in each layer, two residual blocks comprised by two linear layers with SiLu.
We primarily utilized NVIDIA Tesla V100 GPUs with 32 GB of memory.
@@ Appendix C Experimental methodology: Computing the Singular Values of the Jacobian of the Score Function
For computing the singular values, we use the procedure from (Stanczuk et al.,, 2022) reported in algorithm 1. For the linear models and MNIST models we used a symmetrized version which we empirically found to be more stable, reported in algorithm 2.
@@
0:  $s_{\theta}$ trained diffusion model (score), $t_{0}$ sampling time, $K$ number of score vectors.
@@
4:  for $i=1,...,K$ do
5:     Sample $x_{t_{0}}^{(i)}\sim\mathcal{N}(x_{t_{0}}|x_{0},\sigma_{t_{0}}^{2}I)$
6:     Append $s_{\theta}(x_{t_{0}}^{(i)},t_{0})$ as a new column to $S$
7:  end for
8:  $(s_{i})_{i=1}^{d},(v_{i})_{i=1}^{d},(w_{i})_{i=1}^{d}\leftarrow\text{SVD}(S)$
@@
0:  $s_{\theta}$ trained diffusion model (score), $t_{0}$ sampling time, $K$ number of score vectors.
@@
4:  for $i=1,...,K$ do
5:     Sample $x_{t_{0}}^{+(i)}\sim\mathcal{N}(x_{t_{0}}|x_{0},\sigma_{t_{0}}^{2}I)$
6:     Sample $x_{t_{0}}^{-(i)}\sim\mathcal{N}(x_{t_{0}}|x_{0},-\sigma_{t_{0}}^{2}I)$
7:     Append $\frac{s_{\theta}(x_{t_{0}}^{+(i)},t_{0})-s_{\theta}(x_{t_{0}}^{-(i)},t_{0})}{2}$ as a new column to $S$
@@
9:  $(s_{i})_{i=1}^{d},(v_{i})_{i=1}^{d},(w_{i})_{i=1}^{d}\leftarrow\text{SVD}(S)$

### v4

@@ A.2 Double variance scenario
Once the density of the eigenvalues is computed one can perform the same change of variables described in Supp. A.1 for the case of single variance, and obtain the density $\rho_{t}(r)$ for the eigenvalues of $J_{t}$.
Fig. 2 reports the evolution in time of the spectral density, as well as its cumulative function $f=0.75$. The cumulative function has been used to estimate the formation of the intermediate gaps to compare with the experiments for the estimation of the data-manifold dimension.
@@ Appendix B Network training and Model Architecture Details
For the image datasets, we used the diffusion setting in (Ho et al.,, 2020). We use the variance scheduler with $\beta_{\text{min}}=10^{-4}$ and $\beta_{\text{min}}=2\times 10^{-2}$, $T=1000$ time steps, and score model backbone (PixelCNN++ (Salimans et al.,, 2017)). Furthermore, for each of the datasets, we adjusted the partameters to account for the different complexity (see Table 1). For each data-set, the number of used training data-points amounts to the full set of data available.
For the linear models, we used a Variance Exploding continuous score model trained with 2M steps (batch size 128). The model had a Residual architecture with size 128 hidden channels in each layer, two residual blocks comprised by two linear layers with SiLu. In this case, the number of used training-data amounts to $2\cdot 10^{5}$ synthetic examples generated according to Section 6.1.
For all experiments we primarily utilized NVIDIA Tesla V100 GPUs with 32 GB of memory.
@@ Appendix C Experimental methodology: Computing the Singular Values of the Jacobian of the Score Function
For computing the singular values, we used an improved version of the procedure from (Stanczuk et al.,, 2022) that is reported in Algorithm 1. As a difference from the methodology described in the literature, we perturb the position $\boldsymbol{x}_{0}$ where we compute the score function $s_{\theta}$ along strictly orthogonal directions. Such orthogonal perturbations result from a Gaussian sampling of $d$ vectors that are subsequently orthogonalized: these perturbed images are not far from other examples present in the training-set. Choosing orthogonal perturbations instead of random ones significantly reduces the anomalous slope appearing when plotting the ordered singular values that can hide some intermediate gaps (see Fig. $5$ in Stanczuk et al., (2022)). Fig. 9 compares the standard SVD method from the literature with our technique by applying it on model trained through from the CelebA and MNIST data-set.
Figure 9: Comparison between the standard SVD method employed by Stanczuk et al., (2022) (left) and the new orthogonalized method for the visualization of the intermediate gaps (right) for a neural network trained on the CelebA and MNIST data-set as described in Section B. Spectra are measured according to Supp. C.
Moreover, for the linear models and MNIST models we used a symmetrized version which we empirically found to be more stable, reported in algorithm 2.
Once the SVs have been obtained through the algorithms, they will be normalized by the highest value in the set and subsequently ordered in a decreasing fashion. In most cases, the very first singular value is removed from the set, due to a divergent behaviour that is intrinsic to the neural network training.
@@
0:  $s_{\theta}$ trained diffusion model (score), $t_{0}$ sampling time.
@@
4:  for $i=1,...,d$ do
5:     Sample $x_{t_{0}}^{(i)}\sim\mathcal{N}(x_{t_{0}}|x_{0},\sigma_{t_{0}}^{2}I)$ perturbations
6:  end for
7:  $(x_{t_{0}}^{(i)})_{i=1}^{d}\leftarrow(\tilde{x}_{t_{0}}^{(i)})_{i=1}^{d}$ orthogonalized perturbations.
8:  for $i=1,...,d$ do
9:     Append $s_{\theta}(\tilde{x}_{t_{0}}^{(i)},t_{0})$ as a new column to $S$
10:  end for
11:  $(s_{i})_{i=1}^{d},(v_{i})_{i=1}^{d},(w_{i})_{i=1}^{d}\leftarrow\text{SVD}(S)$
@@
0:  $s_{\theta}$ trained diffusion model (score), $t_{0}$ sampling time.
@@
4:  for $i=1,...,d$ do
5:     Sample $x_{t_{0}}^{+(i)}\sim\mathcal{N}(x_{t_{0}}|x_{0},\sigma_{t_{0}}^{2}I)$ right perturbations
6:     Sample $x_{t_{0}}^{-(i)}\sim\mathcal{N}(x_{t_{0}}|x_{0},\sigma_{t_{0}}^{2}I)$
7:     $x_{t_{0}}^{-(i)}\leftarrow 2x_{0}-x_{t_{0}}^{-(i)}$ left perturbations
@@
9:  $(x_{t_{0}}^{+(i)},x_{t_{0}}^{-(i)})_{i=1}^{d}\leftarrow(\tilde{x}_{t_{0}}^{+(i)},\tilde{x}_{t_{0}}^{-(i)})_{i=1}^{d}$ orthogonalized perturbations
10:  for $i=1,...,d$ do
11:     Append $\frac{s_{\theta}(\tilde{x}_{t_{0}}^{+(i)},t_{0})-s_{\theta}(\tilde{x}_{t_{0}}^{-(i)},t_{0})}{2}$ as a new column to $S$
12:  end for
13:  $(s_{i})_{i=1}^{d},(v_{i})_{i=1}^{d},(w_{i})_{i=1}^{d}\leftarrow\text{SVD}(S)$
Appendix D Linear manifold model Hypothesis
The manifold hypothesis is a fundamental concept in machine learning. It states that $d$-dimensional data lie on a lower $m$-dimensional manifold embedded in the high-dimensional space, often called the ambient space (Fefferman et al.,, 2016). Whether it exists, such manifold is supposed to present a local curvature in the ambient space, i.e. it is not representable as a $d$-dimensional hyperplane. The hidden manifold model proposed by Goldt et al., (2020) is able to reproduce data $\boldsymbol{y}^{\mu}$ that are embedded in a latent $m$-dimensional space, as
$\boldsymbol{y}^{\mu}=\Phi\left(F\boldsymbol{z}^{\mu}\right),$ (60)
where $\Phi(\boldsymbol{x})$ is a non-linear function of the $d$-dimensional vector $\boldsymbol{x}$, $F\in\mathbb{R}^{d\times m}$ is a rectangular matrix that projects latent vectors $\boldsymbol{z}^{\mu}$ on the manifold. Nevertheless, in order to apply the replica method from statistical physics and compute the distribution of the eigenvalues of the Jacobian of the score-function, we constrain ourselves to the simpler case of a linear-manifold, i.e. $\Phi(\boldsymbol{x})=\boldsymbol{x}$ (see Section 6.1). Indeed this might sound as a limitation to the reproducibility of our results to more realistic data-sets, e.g. natural images. However we claim that, for ordinary diffusive diffusion models in the variance exploding setup, the theory developed for the linear model still applies to non-linear manifold instances, due to the following reasons:
  1. 1.
    The diffusive trajectories at large $t$, where the trivial phase occurs, are sampled by a probability distribution that is smooth, due to the Gaussian kernel implied by the stochastic process in Eq. 4. As a consequence, the stable latent set defined in Eq. 5 will be approximately linear.
  2. 2.
    The diffusive trajectories at small $t$, where the manifold coverage and manifold consolidation phases occur, explore a region contained into a ball of radius proportional to $\sqrt{t}$, that is supposed to be smaller than then the inverse local curvature of the manifold.
We now compare the distribution of the singular values of the Jacobian obtained from the linear manifold model and the non-linear one. Specifically, we will repeat the plot in Fig. 3 (left and central panels) with a toy data-set generated as $\boldsymbol{y}^{\mu}=F\boldsymbol{\tilde{z}}^{\mu}$, with $\boldsymbol{\tilde{z}}^{\mu}=\boldsymbol{z}^{\mu}/\lVert\boldsymbol{z}^{\mu}\rVert$. As a consequence, the new data will now live on a $(d-1)$-dimensional ellipsoid inscribed in the original $d$-dimensional hyperplane. The results are reported in Fig. 10 and they show no evident discrepancy between the linear and non-linear case, as predicted by our argument. The only small difference consists of a slight delay in time of the gap phenomenology that is present in the non-linear manifold data which impedes the final gap to fully open at the smallest observable time.
Figure 10: Ordered singular value spectrum estimated experimentally from the analysis of a data-set living on a non-linear manifold built according to Supp. D where parameters are chosen as in Fig. 3. The neural network is trained as prescribed in Supp. B and spectra are measured according to Supp. C.
Appendix E Additional Experiments
We provide here additional results regarding the disposition of the ordered singular values of the Jacobian of the score function of models trained according to B and C. Specifically, each plot represents the jacobian measured with respect to one single image in the training data-set. We conclude that all images present the same three-phase phenomenology, as a general feature of the data-sets.
Figure 11: Spectrum of the ordered SVs of the Jacobian for a model trained on MNIST. Each panel is relative to a different data-point in the full set.
Figure 12: Spectrum of the ordered SVs of the Jacobian for a model trained on Cifar10. Each panel is relative to a different data-point in the full set.
Figure 13: Spectrum of the ordered SVs of the Jacobian for a model trained on CelebA. Each panel is relative to a different data-point in the full set.